OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 −x − y = −7 −y + 7 −y − 7 y + 7 y − 7

OpenStudy (anonymous):

-x-y=-7 -x=-7+y x=-y+7

OpenStudy (anonymous):

So what do u think is the answer?

OpenStudy (anonymous):

wow u r amazing thx love u (as a frnd)! @sauravshakya

OpenStudy (anonymous):

can u ans more questions?

OpenStudy (anonymous):

cause i m the worst at math

OpenStudy (anonymous):

I cant....... but I will surley teach u

OpenStudy (anonymous):

k thx @sauravshakya i love u :)

OpenStudy (anonymous):

Opps...... Luv u @tej18

OpenStudy (anonymous):

whats ur name?